### v1

# thanks: These authors contribute equallythanks: These authors contribute equally

Abstract
Physical reservoir computing provides a powerful machine learning paradigm that exploits nonlinear physical dynamics for efficient information processing. By incorporating quantum effects, quantum reservoir computing gains superior potential in machine learning applications, for the quantum dynamics are exponentially costly to simulate classically. Here, we present a novel quantum reservoir computing approach based on correlated quantum spin systems, exploiting natural quantum many-body interactions to generate reservoir dynamics, thereby circumventing the practical challenges of deep quantum circuits. Our experimental implementation supports nontrivial quantum entanglement and exhibits sufficient dynamical complexity for high-performance machine learning. We achieve state-of-the-art performance in experiments on standard time-series benchmarks, reducing prediction error by one to two orders of magnitude compared to previous quantum reservoir experiments. In long-term weather forecasting, our 9-spin quantum reservoir delivers greater prediction accuracy than classical reservoirs with thousands of nodes. This represents a first experimental demonstration of quantum machine learning outperforming large-scale classical models on real-world tasks.
Introduction.—The rapid growth of artificial intelligence has highlighted the demand for novel computational paradigms capable of real-time processing of temporally structured, data-intensive information streams. Physical reservoir computing (PRC) offers a promising solution by harnessing the intrinsic dynamics of physical substrates as computational resources within a machine learning framework Tanaka et al. (2019); Nakajima (2020). In this approach, inputs are projected into the high-dimensional space of a fixed dynamical system that simultaneously performs processing and memory, thus avoiding the von Neumann bottleneck. The output is obtained by only training a linear readout layer, enabling fast inference with minimal learning cost. PRC can be implemented in a broad range of physical systems that provide a high-dimensional state space, nonlinear input–output mapping, and fading memory Konkoli (2017), including memristors Moon et al. (2019); Milano et al. (2022); Liu et al. (2022); Liang et al. (2024), spintronics Torrejon et al. (2017); Romera et al. (2018); Grollier et al. (2020); Lee et al. (2024), and optical systems Brunner et al. (2013); Vandoorne et al. (2014); Larger et al. (2017); Rafayelyan et al. (2020). As larger and more complex dynamical systems tend to enhance computational capacity Kaplan et al. (2020); Raghu et al. (2017); Lukoševičius and Jaeger (2009), this framework naturally extends into the quantum regime. Quantum reservoir computing (QRC) leverages the exponentially large Hilbert space and complex correlations of quantum systems to offer unique advantages in solving challenging tasks. Given that some noisy intermediate-scale quantum (NISQ) systems have already outperformed classical simulators Arute et al. (2019); Zhong et al. (2020); Madsen et al. (2022), current quantum devices hold strong potential for realizing QRC with computational capacity and expressiveness beyond classical models.
Despite its strong potential, QRC has so far been explored primarily through theoretical studies and numerical simulations on various idealized quantum models, while experimental realizations remain relatively limited Mujal et al. (2021); Ghosh et al. (2021). Most implementations rely on deep quantum circuits to generate the complex dynamics required for QRC Hu et al. (2024); Pfeffer et al. (2022); Chen et al. (2020); Kubota et al. (2023); Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). However, such circuits often demand substantial depth, making them vulnerable to gate errors and dissipation, thereby limiting the fidelity and performance of the QRC in the experiments. Furthermore, while quantum systems inherently possess exponentially large Hilbert spaces, current readout schemes are typically limited to local measurements, restricting the full exploitation of the high-dimensional nature of the quantum dynamics.
In this Letter, we present an experimental realization of quantum reservoir computing based on a quantum spin system. To address the abovementioned challenges, we introduce two key innovations: the integration of both coherent spin-spin interactions and incoherent spin relaxation processes to enrich dynamical complexity of the reservoir, and the use of time-multiplexed measurements to increase the effective readout capacity of the reservoir Fujii and Nakajima (2017). Building on these, we achieve the state-of-the-art performance in QRC experiments on the widely used NARMA benchmark task, reducing the prediction error by one to two orders of magnitude compared to previous results Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). We apply the QRC protocol to weather forecasting, which is crucial for countless social and economic activities Price et al. (2025), yet challenging due to the chaotic nature of the atmosphere Kalnay (2003); Palmer (1993). In long-term predictions, our 9-spin quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes. Our results suggest practical quantum advantages in time-series prediction may be achievable with current quantum hardware. To our knowledge, this represents a first experimental demonstration of a quantum machine learning system that outperforms large-scale classical networks on realistic datasets, where previous advantages were typically shown via numerical simulations or limited to synthetic datasets Biamonte et al. (2017); Huang et al. (2021); Liu et al. (2021); Cerezo et al. (2022); Caro et al. (2022); Jäger and Krems (2023); Xia et al. (2023); Kornjača et al. (2024).
Figure 1: Schematic diagram of quantum reservoir computing. The input series $\{s_{k}\}$ is mapped to the control parameters $\{\theta_{k}\}$ of a sequence of quantum operations, which are sequentially applied to a quantum reservoir comprising a spin network. The reservoir state is probed through measurements of partial observables, producing a readout vector with components $x^{i}_{k}=\langle O^{i}\rangle_{k}={\rm Tr}(\rho_{k}O^{i})$. The final output is computed as a linear combination, $\hat{y}_{k}={\bf x}_{k}^{\top}{\bf w}$.
Model design.—We introduce the general framework of QRC for temporal information processing, as illustrated in Fig. 1. Consider a pair of discrete scalar sequences: the input $\{s_{k}\}$ and the target output $\{y_{k}\}$, where $y_{k}$ depends on a finite input history up to time step $k$, denoted by $y_{k}=f(s_{k},\ldots,s_{k-\Delta})$. The goal of QRC is to approximate this relationship by generating an output sequence $\{\hat{y}_{k}\}$ that closely matches the target. The inputs $\{s_{k}\}$ are encoded into a control sequence and sequentially applied to an $N$-qubit quantum reservoir, mapping the information into a $4^{N}$-dimensional Hilbert space. Between successive inputs, the reservoir undergoes a fixed intrinsic evolution, which integrates new inputs with the memory of past inputs. To extract the information, specific observables are measured at each time step, yielding a sequence of readout vectors ${\bf x}_{k}$} with $x_{k}^{i}={\rm Tr}(\rho_{k}O^{i})$. The output is computed as a weighted sum of readout components, $\hat{y}_{k}={\bf x}_{k}^{\top}{\bf w}$, with a bias term incorporated by augmenting ${\bf x}_{k}$ with $x^{0}_{k}=1$. Following standard machine learning practices, the readout weights $\bf w$} are optimized on a training dataset by minimizing the error $\sum_{k=1}^{L}\left(y_{k}-\hat{y}_{k}\right)^{2}$, and the performance of the resulting QRC is then evaluated on test data.
Figure 2: Schematic of experimental setup and readout schemes. (a) Circuit diagram of the QRC experimental implementation. The reservoir, composed of carbon and proton spins, undergoes repeated cycles of input-dependent rotations $\{\phi_{k},\theta_{k}\}$ followed by intrinsic evolution under $\exp(\tau\mathcal{L}_{H})$. The experiment adopts a rewinding protocol, where each experiment at time step $k$ is restarted from cycle $k-n_{\rm wo}$ to wash out initial state dependence, and readout is performed on the proton spins at the end of the circuit. Two readout schemes are considered: (b) Single-time readout: analogous to capturing multiple images of a static system by artificially changing the viewpoint. A fixed set of observables $\{O^{i}\}$ (e.g. Pauli operators) is measured on the reservoir state $\rho_{k}$, yielding the readout vector ${\bf x}_{k}=[\langle O^{1}\rangle_{k},\langle O^{2}\rangle_{k},\dots]^{\top}$ where $\langle O\rangle_{k}={\rm Tr}(\rho_{k}O)$. (c) Time-multiplexed readout: analogous to filming a moving system from a fixed viewpoint, where other aspects naturally rotate into view. A single observable $O$ is sampled at multiple time points during the reservoir’s evolution, equivalently measuring the Heisenberg‑picture operators $\{O(t_{i})\}$, yielding the readout vector ${\bf x}_{k}=[\langle O(t_{1})\rangle_{k},\langle O(t_{2})\rangle_{k},\dots]^{\top}$.
The computational capabilities of QRC essentially depend on both the dynamics of the quantum reservoir and the number of readout components Mart´ınez-Peña et al. (2021); Xia et al. (2022); Dambre et al. (2012); Mart´ınez-Peña et al. (2023). Generally, QRC dynamics consist of unitary and non-unitary components. The unitary evolution, realized via analog Hamiltonian dynamics or digital quantum circuits, transforms and spreads information throughout the system while preserving coherence and generating entanglement. In addition to enriching the dynamics, the non-unitary part is essential for retaining memory of recent inputs by gradually diminishing the influence of past states, a property known as fading memory Boyd and Chua (1985). Previous protocols have achieved this through qubit resetting Fujii and Nakajima (2017); Hu et al. (2024) or engineered dissipative channels Chen et al. (2020); Sannia et al. (2024), but these methods often require intricate control techniques or additional experimental resources.
Here, we adopt a more efficient method by leveraging the inherent relaxation, which commonly exists in natural quantum systems. Thus, the system evolution is governed by the Lindbladian superoperator $\dot{\rho}=\mathcal{L}_{H}[\rho]=-i\left[H,\rho\right]+\mathcal{R}[\rho]$, where $H$ is the system’s Hamiltonian and $\mathcal{R}$ denotes the relaxation channel. The overall state update is then expressed as
$\rho_{k}=e^{\tau\mathcal{L}_{H}}\circ\mathcal{U}_{s_{k}}\left[\rho_{k-1}\right],$ (1)
where $e^{\tau\mathcal{L}_{H}}$ denotes the propagator for evolution over duration $\tau$, and $\mathcal{U}_{s_{k}}[\rho]=U_{s_{k}}\rho U^{\dagger}_{s_{k}}$ represents the unitary channel used to encode the classical input $s_{k}$. In spin systems, two dominant relaxation processes are longitudinal relaxation (with time constant $T_{1}$), which drives the system toward thermal equilibrium via energy dissipation, and transverse relaxation (with time constant $T_{2}$), which causes coherence decay through phase randomization Levitt (2008). These processes correspond to generalized amplitude damping and dephasing channels, respectively Nielsen and Chuang (2001). Our results indicate that $T_{1}$ relaxation plays a pivotal role in ensuring QRC performance (see Supplementary Material sup ), in agreement with recent numerical studies Domingo et al. (2023); Kubota et al. (2023); Monzani et al. (2024). By incorporating inevitable relaxation as a computational resource, our method overcomes coherence time limitations and eliminates the need for qubit resets, enabling prolonged experimental temporal processing with minimal control.
Despite the rich dynamics of quantum reservoirs, the information processing capacity of a dynamical system is fundamentally limited by the number of independent readout functions Dambre et al. (2012). While quantum systems provide access to an exponentially large state space, extracting complete information through state tomography becomes increasingly challenging for large systems, particularly when high-order correlations are involved Nielsen and Chuang (2001). Consequently, most existing schemes focus on extracting local observable information like $\langle\sigma_{z}^{i}\rangle$ Chen et al. (2020); Suzuki et al. (2022); Kubota et al. (2023); Yasuda et al. (2023); Monzani et al. (2024), which capture only a small subset of the reservoir’s degrees of freedom, thus failing to fully exploit its computational potential.
To address this limitation, we employ time multiplexing, a technique that effectively increases the number of readouts by performing repeated measurements at regular intervals during the reservoir’s evolution Fujii and Nakajima (2017), as illustrated in Fig. 2(c). When input encoding and evolution are decoupled in QRC, time multiplexing can be viewed as a form of temporal tomography, where fixed observables are measured and information about other components is continuously mapped onto them through evolution Yang et al. (2020); Chantasri et al. (2020); Ghosh et al. (2020). In our implementation, we utilize the free induction decay signal, a natural manifestation of time multiplexing in quantum systems, to capture temporal information, with experimental results showing significant performance improvement.
High-accuracy temporal prediction.—To demonstrate the performance of our approach, we perform experimental benchmarks on a nuclear magnetic resonance (NMR) quantum platform. The initial task we consider is the nonlinear autoregressive moving average (NARMA) simulation, a widely adopted benchmark for temporal processing capabilities Hochreiter (1997); Atiya and Parlos (2000). In accordance with previous work Fujii and Nakajima (2017); Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024), the input sequence is constructed as a superposition of sine waves, while the target sequence is generated by specific recurrent dynamics, for example:
$y_{k+1}=\alpha y_{k}+\beta y_{k}\left(\sum_{i=0}^{n-1}y_{k-i}\right)+\gamma s_{k-n+1}s_{k}+\delta,$ (2)
where $n$ determines the number of past time steps influencing the current output, with the task denoted by NARMAnn. Here, we allocate 400 steps for training and 100 steps for testing. Readout weights are optimized via ridge regression Hoerl and Kennard (1970): ${\bf w}^{*}=\left({\bf X}^{\top}{\bf X}+\lambda{\bf I}\right)^{-1}{\bf X}^{\top}{\bf y}$, where ${\bf X}=\left[{\bf x}_{1},\ldots,{\bf x}_{400}\right]^{\top}$ and ${\bf y}=\left[y_{1},\ldots,y_{400}\right]^{\top}$ denote the readout matrix and target vector for training, and $\lambda$ is the regularization strength. By fitting separate $\bf w$$^{*}$ for each $n$ on the same $\bf X$}, the reservoir simultaneously emulates multiple NARMA systems, a capability commonly referred to as multitasking Fujii and Nakajima (2017).
Fig. 2(a) presents the circuit schematic for our QRC experiments. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule. The spin reservoir undergoes successive cycles of input encoding and free evolution according to Eq. (1). For NARMA task, its scalar input is normalized to $\bar{s}_{k}\in[0,1]$ and encoded via an $x$-axis radio-frequency pulse on proton nuclei, effecting a global rotation $U_{s_{k}}=\prod_{n=5}^{9}R^{(n)}_{x}\left(\arcsin(\bar{s}_{k})\right)$. The system Hamiltonian is
$H=\sum_{i=1}^{9}\pi\nu_{i}\sigma^{i}_{z}+\sum_{1\leq i<j\leq 9}\frac{\pi}{2}J_{ij}\sigma^{i}_{z}\sigma^{j}_{z},$ (3)
where $\nu_{i}$ represents the chemical shift of spin-$i$ and $J_{ij}$ denotes the coupling between spin-$i$ and spin-$j$ sup . To circumvent the measurement back-action, we adopt the rewinding protocol Mujal et al. (2023), where the experiment of time step $k$ is performed with input cycles from $k-n_{\rm wo}$ to $k$. The first $n_{\rm wo}$ cycles serve to wash out the initial state dependence under the premise of fading memory, thus allowing the experiment to start from any state without specific preparation. The washout length $n_{\rm wo}$ depends on the system’s relaxation rates and evolution durations, and is determined through preliminary experiments sup .
After the $n_{\rm wo}+1$ cycles, readout is performed on the proton spins using two schemes, as illustrated in Fig. 2(b) and (c). Single‑time readout refers to the conventional measurement of a fixed set of observables, here chosen as single-qubit Pauli operators. As $\rm H_{3}$$\rm H_{5}$ are indistinguishable methyl protons, only nine distinct expectation values, $\langle\sigma^{m}_{\alpha}\rangle$ ($\alpha=x,y,z$, $m={\rm H_{1}},{\rm H_{2}},{\rm methyl}$), are accessible. Time-multiplexed readout is realized using the free induction decay (FID) signal, which uniformly samples the nuclear transverse magnetization during free evolution Levitt (2008):
$S(t)\propto{\rm Tr}\left\{e^{t\mathcal{L}_{H}}\left[U\rho U^{\dagger}\right]\,\sum_{m=5}^{9}\left(\sigma^{m}_{y}+i\sigma^{m}_{x}\right)\right\}.$ (4)
In the experiments, we apply a $\pi/2$ pulse before FID readout ($U=\prod_{m=5}^{9}R^{m}_{x}(\pi/2)$) and apply a Fourier transform to the FID signal to isolate the reservoir response and reduce dimensionality, yielding 653 spectral features. Further details are provided in Supplementary Material sup .
Figure 3: High accuracy temporal prediction with NARMA benchmark. (a) Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, using quantum reservoirs with different configurations. Red and blue lines represent evolution durations of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time multiplexing (TM) and single-time readout schemes. (b) Prediction results for NARMA tasks with $n=10$ and $20$, utilizing time-multiplexed readout with $\tau=$0.01\text{\,}\mathrm{s}$\$.
$\tau=$0.3\text{\,}\mathrm{s}$\$ $\tau=$0.01\text{\,}\mathrm{s}$\$
NMSE TM TM
NARMA2 $2.76\times 10^{-6}$ $\bf 1.74\times 10^{-7}$ $1.03\times 10^{-5}$ $2.10\times 10^{-7}$
NARMA5 $5.03\times 10^{-4}$ $4.44\times 10^{-5}$ $2.01\times 10^{-3}$ $\bf 3.44\times 10^{-5}$
NARMA10 $1.16\times 10^{-3}$ $3.39\times 10^{-4}$ $1.73\times 10^{-3}$ $\bf 5.84\times 10^{-5}$
NARMA15 $1.78\times 10^{-3}$ $5.64\times 10^{-4}$ $1.96\times 10^{-3}$ $\bf 6.37\times 10^{-5}$
NARMA20 $1.66\times 10^{-3}$ $7.15\times 10^{-4}$ $8.68\times 10^{-4}$ $\bf 4.34\times 10^{-5}$
Table 1: List of normalized mean squared errors ($\mathrm{NMSE}=\sum_{k=1}^{L}\left(y_{k}-\hat{y}_{k}\right)^{2}\big{/}\sum_{k=1}^{L}y_{k}^{2}$) for NARMA tasks with orders $n=2$, 5, 10, 15, and 20, using QRC under various configurations. The minimum error in each task is highlighted.
We evaluate quantum reservoirs with different evolution durations and readout schemes on NARMA tasks of orders $n=2$ to 20. The results are shown in Fig. 3(a), with performance quantified by the coefficient of determination,
$R^{2}=1-\frac{\sum_{k=1}^{L}\left(y_{k}-\hat{y}_{k}\right)^{2}}{\sum_{k=1}^{L}\left(y_{k}-\bar{y}\right)^{2}},$ (5)
where $\bar{y}=\frac{1}{L}\sum_{k=1}^{L}y_{k}$. For direct comparison with prior work, we also report the corresponding normalized mean square errors in TABLE 1; our best cases achieve error reductions of 1-2 orders of magnitude relative to recent experimental benchmarks Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). Time multiplexing markedly improves performance over single-time readout, especially for $\tau=$0.01\text{\,}\mathrm{s}$\$, where predictions are nearly perfect across all NARMA orders (see NARMA10 and 20 in Fig. 3(b) and additional results in Supplementary Material sup ). As the relaxation effects become more pronounced with longer evolution times, the information from high-order correlations is significantly diminished at $\tau=$0.3\text{\,}\mathrm{s}$\$, thereby reducing the benefit of time multiplexing. In the Supplementary Material, it is shown that the quantum reservoir significantly outperformed its classical spin-based counterpart, which yields an average $R^{2}\textless 0.67$ across NARMA2 to NARMA20, highlighting the role of quantum correlations in computation sup .
Figure 4: Weather forecasting via quantum reservoir computing. (a) Experimental results for single-step-ahead forecasting of temperature and humidity using quantum reservoir computing. The black and red lines denote the target and predicted trajectories, respectively. (b) Multi-step-ahead forecasting performance comparison between echo state networks (ESNs) with varying node counts and QRC methods. $R^{2}$ is averaged over 100 random realizations for each ESN($m$), with shaded regions representing the standard deviation as error bars.
Realistic weather forecasting.—To show the practical utility of our approach, we conduct weather forecasting on a daily climate dataset from Delhi, a benchmark commonly employed in the machine learning community dai (2019). Our analysis focuses on predicting the multivariate time series of temperature and humidity, ${\bf y}_{k}=\left[{\rm Temp}(k),{\rm Humi}(k)\right]$. We first perform single-step-ahead forecasting, where ${\bf y}_{k}$ serves as the $k$-th input and ${\bf y}_{k+1}$ as the corresponding target. The dataset is divided into 374 washout steps, 600 training steps, and 600 testing steps. The normalized temperature and humidity data are encoded as angles $\theta_{k}$ and $\phi_{k}$ for global rotations applied to the proton and carbon nuclei, respectively, with an input interval of $\tau=$0.03\text{\,}\mathrm{s}$\$. Fig. 4(a) presents the single-step forecasting results alongside the target values from the test set. Real weather data exhibit substantial stochastic fluctuations, which pose challenges for precise forecasting. Nevertheless, QRC predictions successfully capture the overall trends and demonstrate close agreement with local variations in several segments of the time series.
We subsequently assess multi-step-ahead forecasting, in which predictability is constrained by the rapid error growth inherent to the chaotic atmospheric system Kalnay (2003); Palmer (1993). Here, a direct multi-step forecast strategy is employed Taieb et al. (2012), wherein a separate model is trained for each forecast horizon $h$: ${\bf y}_{k+h}=f_{h}({\bf y}_{k},{\bf y}_{k-1},{\bf y}_{k-2},\ldots)$. This is implemented through multitasking in QRC, where distinct readout weights are trained for each target sequence $\{{\bf y}_{k+h}\}$ while utilizing the same input sequence $\{{\bf y}_{k}\}$. To evaluate the potential quantum advantages, we compare the performance of QRC with that of echo state networks (ESNs) Jaeger and Haas (2004), with results displayed in Fig. 4(b). ESNs with 500, 1000, 5000, and 10000 nodes (denoted as ESN($m$) for $m$-node networks) are evaluated, showing improved performance with increasing size until saturation, where the curves of ESN(5000) and ESN(10000) nearly overlap, indicating diminishing returns. For temperature forecasting, QRC attains accuracy comparable to that of ESN(1000), and even exceeds its average at larger forecast horizons $h$. Moreover, incorporating nonlinear post-processing may further enhances QRC’s performance: support vector regression (SVR) with a radial basis function (RBF) kernel Awad et al. (2015) enables QRC to achieve higher accuracy than ESN(10000), whereas ESNs gain no similar improvementsup . We attribute QRC’s strong performance to the superior expressibility Sim et al. (2019) and unique information processing capabilities Kubota et al. (2023) inherent to quantum models. Although humidity prediction is inherently more challenging due to its environmental sensitivity, QRC still achieves performance comparable to that of ESN(10000). These results show the capability of quantum systems to address complex dynamical prediction tasks, highlighting the potential quantum advantages in real-world applications with current NISQ devices.
Conclusion.—Building on the natural dynamics of a spin system, we develop a hardware-efficient quantum reservoir computing scheme that enables accurate time-series predictions in practical machine learning tasks. Implemented on a 9-spin system with time-multiplexed readouts, our approach achieves one to two orders of magnitude lower error compared to previous QRC experiments on NARMA benchmarks. In long-term weather forecasting, our quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes, suggesting that practical quantum advantages in time-series prediction may be attainable with current quantum hardware. We anticipate the computational capacity of our QRC protocol can be further improved by increasing the complexity of the reservoir dynamics, for example through pulse-engineered refinement of internal interactions, or by scaling to larger quantum processors. Moreover, integrating QRC with many-body phenomena, including dynamical phase transitions Mart´ınez-Peña et al. (2021); Xia et al. (2022); Kobayashi and Motome (2023, 2025), quantum scar states Serbyn et al. (2021); Bravo et al. (2022), and time crystals Zhang et al. (2017), represents a promising direction for future research, for it could help uncover physical principles to improve QRC performance systematically.
Acknowledgements.
This work is supported by Innovation Program for Quantum Science and Technology (Grants No. 2024ZD0302000, 2021ZD0303205, 2024ZD0300100), National Natural Science Foundation of China (Grants No. T2388102, 12261160569, 92165108), National Basic Research Program of China (Grants No. 2021YFA1400900), Shanghai Municipal Science and Technology (2019SHZDZX01, 24DP2600100), and the XPLORER Prize. The authors acknowledge helpful discussion with the Anhui Meteorological Bureau of China.
Supplemental Material for High-Accuracy Temporal Prediction via Experimental Quantum Reservoir Computing in Correlated Spins
.1 I.    Experimental setup
In this section we present the experimental setup employed in this work. Our experiments were conducted using a Bruker Avance III 400 MHz spectrometer at 298 K. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule, the structure of which is shown in TABLE 2. The methyl protons $\rm H_{3}$, $\rm H_{4}$ and $\rm H_{5}$ are chemically and magnetically equivalent and therefore indistinguishable. TABLE 2 lists the chemical shifts $\nu_{i}$ (diagonal elements) and the $J$-coupling constants $J_{ij}$ (off-diagonal elements), along with the corresponding relaxation times. In particular, the $T_{2}$ times account for dephasing induced by magnetic field inhomogeneities (referred to as $T_{2}^{*}$ in NMR Levitt (2008)).
Table 2: Sample parameters and molecular structure of nine-spin crotonic acid sample. The left portion presents chemical shifts (diagonal elements) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times.
The preprocessed input component is encoded in the pulse area of a radiofrequency (RF) field, with its phase aligned along the $x$-axis and frequency resonant with designated nuclei (in this work we consider inputs with ${\rm dim}\leq 2$, while higher‑dimensional inputs can be encoded using general parameterized circuits Li et al. (2022); Caro et al. (2021)). Each pulse, with a duration of approximately $10^{-5}$ s, is short enough that the system’s free evolution during the pulse can be neglected, thereby effectively implementing a global rotation on all proton spins. Although slight deviations in the rotation angle and axis may arise from detuning between the chemical shift and the RF frequency, this effect remains fixed and is incorporated within the quantum reservoir. The pulses are applied at regular intervals $\tau$, during which the system evolves under its natural Hamiltonian and relaxation dynamics without external control. The readout is performed at the end of the evolution period to measure the protons, since the signal-to-noise ratio for carbon is significantly lower. Consequently, the carbon spins are treated as an inaccessible bath, demonstrating that the computational power of the system can be harnessed solely by accessing a subset of its spins. Two types of readout schemes are implemented: one scheme involves measuring the observable expectations $\langle\sigma^{m}_{x,y,z}\rangle$ with $m={\rm H_{1}},{\rm H_{2}}$, and the methyl protons, while the time multiplexing readout scheme is described in the subsequent section. The readout vector dataset is normalized using Z-score standardization followed by ridge regression, with the regularization strength determined via 10-fold cross validation.
In our experiments, the number of washout steps $n_{\rm wo}$ was determined by extending the washout sequence until the NMR spectral readout (see FIG. 5(b)) remained invariant across different initial states. For NARMA tasks, we used $n_{\rm wo}=100$ for $\tau=$0.3\text{\,}\mathrm{s}$\$ and 1000 for $\tau=$0.01\text{\,}\mathrm{s}$\$, which may be larger than the minimum required to achieve convergence. As described in the main text, a rewinding protocol was employed in our experiments Mujal et al. (2023); that is, each experiment is reinitialized with last $n_{\rm wo}$ input-evolution cycles (from $k-n_{\rm wo}$ to k1k-1) to reprepare the reservoir to $\rho_{k-1}$, then perform the input-evolution-readout procedure for the $k$-th step. Other experimental protocols have been reported in recent years Mujal et al. (2023); Kobayashi et al. (2024), such as the weak measurement protocol that enables online data processing without restarting experiments, which we plan to explore in future studies.
.2 II.    FID signal and NMR spectrum as readout
[…]
Here we outline the mathematical details. Assume that a unitary transformation $U$ is applied to the final state $\rho$ to be measured, and the FID signal is subsequently recorded from the first qubit. This signal can be expressed as (with $\omega_{i}=2\pi\nu_{i}$):
[…]
Figure 5: (a) The real part of the free induction decay (FID) signal obtained from the thermally equilibrated proton spins of crotonic acid, truncated to show the first 4096 data points (half of the total acquisition time), with the signal amplitude enhanced by the spectrometer’s receiver amplification. (b) The real part of corresponding NMR spectrum, produced by Fourier transforming the FID signal. The intense peak near 1880 Hz originates from $\rm H_{2}O$ in the solvent. The red, green, and orange boxes highlight the spectral regions of crotonic acid’s $\rm H_{1}$, $\rm H_{2}$, and methyl protons $\rm H_{Me}$ (i.e. the equivalent $\rm H_{3},H_{4},H_{5}$), respectively; the insets display magnified views of these regions. Although the signal intensities of the crotonic acid protons are two to three orders of magnitude lower than that of the water peak, they still exhibit sufficiently high signal-to-noise ratios.
Thus far, we have explicitly detailed the mechanism for extracting state information via time-multiplexing (i.e., the FID signal) under a liquid NMR Hamiltonian, while the more general process is to learn an implicit linear transformation between the temporal measurements and the state representation Yang et al. (2020); Chantasri et al. (2020); Xiao et al. (2024). However, in practice, we use the spectral peaks of the FID signal after Fourier transformation as readout features, as highlighted in FIG. 5(b). This choice is motivated by the fact that the FID signal contains a significant contribution from $\rm H_{2}O$ in the solvent, which produces a dominant peak near 1880 Hz. Moreover, removing signal-independent components (e.g. random noise) can reduce computational complexity and mitigate the risk of overfitting. The spectral peaks can be described analytically as:
$S(\omega)=\int_{0}^{\infty}S(t)e^{i\omega t}{\rm d}t\propto\sum_{\bf q}M_{\bf q}\left[\frac{\lambda}{\left(\omega-\Omega_{\bf q}\right)^{2}+\lambda^{2}}+i\frac{\omega-\Omega_{\bf q}}{\left(\omega-\Omega_{\bf q}\right)^{2}+\lambda^{2}}\right]=\sum_{\bf q}M_{\bf q}\left[A_{\bf q}(\omega)+iD_{\bf q}(\omega)\right],$ (7)
where $A_{\mathbf{q}}(\omega)$ and $D_{\mathbf{q}}(\omega)$ correspond to the absorption and dispersion Lorentzian line shapes, respectively, centered at $\Omega_{\bf q}$ with a linewidth of $\lambda$. Typically, one would extract each component by fitting the spectrum; however, for samples with many spins, overlapping peaks (i.e. when $|\Omega_{\bf q}-\Omega_{\bf q^{\prime}}|\lesssim\lambda$) complicate such fitting. Only single-qubit observable expectations can be obtained by integrating over the corresponding spectral region. Our method, by contrast, directly utilizes the NMR spectral peaks as readouts without explicit post-processing, instead automatically extracting the necessary state information via linear regression. The effectiveness of this approach has been validated through both simulations (Section .5) and experiments (main text). Notably, no classical nonlinear enhancement techniques were introduced in this process. In our implementation, the truncated spectral peaks are merged into 653-dimensional features. While further dimensionality reduction is possible, given the observation that adjacent points on a Lorentzian peak exhibit high linear dependence, such reduction would entail a trade-off between performance and simplicity that we have not explored in detail.
.3 III.    Numerical characterization of spin-based quantum reservoir dynamics
In this section, we presents the numerical implementation of the input injection and the subsequent system evolution in the quantum reservoir. The normalized input $\bar{s}_{k}\in[0,1]$ is applied via a pulse on a specific nuclear channel (e.g., proton), which is equivalent to performing global rotations on the corresponding spins:
$U_{s_{k}}=\prod_{n\in S_{\rm in}}R^{(n)}_{x}\bigl{(}\arcsin(\bar{s}_{k})\bigr{)},$ (8)
where $S_{\rm in}$ denotes the set of qubits constituting the input channel.
The evolution of the system is governed by the Lindblad master equation:
[…]
with the Hamiltonian
$H=\sum_{i=1}^{N}\pi\nu_{i}\,\sigma^{i}_{z}+\sum_{1\leq i<j\leq N}\frac{\pi}{2}J_{ij}\,\sigma^{i}_{z}\sigma^{j}_{z}.$ (10)
We next characterize the relaxation channel $\mathcal{R}$ in detail.
For a single spin, the relaxation channel is given by
$\mathcal{R}[\rho]=\frac{1}{2}\sum_{a=+,-,z}\left(2L_{a}\rho L_{a}^{\dagger}-L_{a}^{\dagger}L_{a}\rho-\rho L_{a}^{\dagger}L_{a}\right),$ (11)
with the collapse operators defined as
$L_{+}=\sqrt{\frac{1-p}{2T_{1}}}\,\sigma_{+},~L_{-}=\sqrt{\frac{1+p}{2T_{1}}}\,\sigma_{-},~L_{z}=\sqrt{\frac{1}{2T_{\phi}}}\,\sigma_{z}.$ (12)
Here, $|\sigma_{+}=\outerproduct{1}{0}$ and $|\sigma_{-}=\outerproduct{0}{1}$ denote the spin raising and lowering operators, respectively. The operators $L_{+}$ and $L_{-}$ define a generalized amplitude damping channel that drives the spin toward its equilibrium state,
$\rho_{\rm eq}=\frac{1}{2}\begin{bmatrix}1+p&0\\ 0&1-p\end{bmatrix}$, (13)
at a rate $T_{1}^{-1}$, where $p$ represents the spin polarization (with $p=1$ recovering the standard amplitude damping channel) Nielsen and Chuang (2001). The operator $L_{z}$ accounts for pure dephasing due to spin-spin interactions and field inhomogeneities, with a characteristic time $T_{\phi}$. Combined with amplitude damping, the effective decoherence rate is given by
$T_{2}^{-1}=\left(2T_{1}\right)^{-1}+T_{\phi}^{-1}.$ (14)
Consequently, the transverse spin polarization decays exponentially as $e^{-t/T_{2}}$ while precessing in the $xy$-plane, whereas the longitudinal polarization along the $z$-axis relaxes toward its equilibrium value according to $p_{z}(t)=p+\left[p_{z}(0)-p\right]e^{-t/T_{1}}$.
For an $N$-spin system, the relaxation dynamics become more complex. In our simulations, we assume that the spins relax independently so that the total relaxation channel is given by
$\mathcal{R}[\rho]=\sum_{n=1}^{N}\mathcal{R}^{(n)}[\rho],$ (15)
with each $\mathcal{R}^{(n)}[\rho]$ characterized by its individual relaxation parameters $T^{(n)}_{1}$ and $T^{(n)}_{2}$.
.4 IV.    The NARMA task setup and experimental prediction results
The NARMA task setup follows the pioneering work on quantum reservoir computing Fujii and Nakajima (2017), which has also been adopted in several subsequent studies Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). Two types of NARMA dynamics are considered:
[…]
where the order $n$ controls the dependence on long time lags. The input sequence is constructed using a superimposed sine wave given by
$s_{k}=0.1\left[\sin\left(\frac{2\pi\bar{\alpha}k}{T}\right)\sin\left(\frac{2\pi\bar{\beta}k}{T}\right)\sin\left(\frac{2\pi\bar{\gamma}k}{T}\right)+1\right],$ (17)
where $(\bar{\alpha},\bar{\beta},\bar{\gamma},T)=(2.11,3.73,4.11,100)$.
FIG. 6 presents the experimental prediction results for NARMA tasks with order $n=2,5,10,15$ and 20, utilizing quantum reservoirs with different configurations. The left and right panels show results with single-time readout ($\langle\sigma^{m}_{x,y,z}\rangle$) and time-multiplexed readout, respectively, while red and blue lines represent evolution durations of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively.
Figure 6: Experimental prediction results for NARMA tasks with $n=2,5,10,15,20$, utilizing quantum reservoirs with different evolution durations and readout schemes.
.5 V.    Numerical simulation of 9-spin reservoir computing for NARMA tasks
To validate the reliability of our experimental results, we performed numerical simulations of the nine-spin quantum reservoir computing model on the NARMA task, following the procedure described in Section .3 and using the parameters listed in Table 2. The simulations were implemented using the Python package QuTiP Lambert et al. (2024). For the polarization setting, because the realistic polarization values in room‑temperature NMR (on the order of $10^{-5}$) can lead to numerical instability when solving the Lindblad equation, we scale the carbon and proton polarizations to $1,\,3.98]\times 10^{-4}$ (proportional to their gyromagnetic ratios) to balance numerical robustness with experimental realism.
The simulation results, presented in FIG. 7(a), qualitatively resemble those in Fig. 3 of the main text. Quantitative differences are attributed to deviations between the simulated dynamics and the actual experimental processes, including relaxation mechanisms that extend beyond independent single-spin relaxation and unidentified systematic errors. In our simulations, we incorporated off-resonance effects during the input stage (i.e., imperfect $R_{x}$ pulses) and introduced Gaussian noise during the readout stage, based on an experimental signal-to-noise ratio (SNR) of approximately 300 in the thermally equilibrated spectrum. Spatial field inhomogeneity was also taken into account; however, it did not significantly affect the results and appears to be incorporated as a fixed mechanism of the reservoir. Accurate modeling of open quantum many-body dynamics remains a significant challenge and warrants further investigation.
Figure 7: Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, obtained from numerical simulations of (a) quantum reservoir computing and (b) classical spin reservoir computing. Red and blue lines represent input intervals of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time multiplexing and single-time measurements.
To elucidate the effect of quantumness, we compare our QRC model with the corresponding classical spin reservoir computing model, obtained by treating the spins as classical unit vectors Kornjača et al. (2024). The dynamics of the $i$-th spin are described by
$\begin{aligned} \frac{{\rm d}S_{x}^{(i)}}{{\rm d}t}&=-\Omega_{i}S_{y}^{(i)}-\frac{S_{x}^{(i)}}{T_{2}^{(i)}}\\ \frac{{\rm d}S_{y}^{(i)}}{{\rm d}t}&=\Omega_{i}S_{x}^{(i)}-\frac{S_{y}^{(i)}}{T_{2}^{(i)}}\\ \frac{{\rm d}S_{z}^{(i)}}{{\rm d}t}&=\frac{1-S_{z}^{(i)}}{T_{1}^{(i)}}\end{aligned},$ (18)
where $\Omega_{i}=2\pi\nu_{i}+2\pi\sum_{j\neq i}J_{ij}S_{z}^{(j)}$ represents the effective instantaneous field exerted on spin-$i$ by the surrounding spins. Hence, the dynamics of $N=9$ classical spins are characterized by $3N=27$ differential equations. The simulation results for the classical spin reservoir, with experimental settings identical to those used for QRC, are shown in FIG. 7(b). These results demonstrate that the optimal performance achievable by the classical spin reservoir remains inferior to that of the quantum reservoir, even when time multiplexing is employed. This is because classical spins lack the additional degrees of freedom, such as quantum correlations, that can be exploited via time multiplexing, thereby highlighting the advantage of quantumness. Moreover, we attribute the improvement achieved by time multiplexing in classical systems to the mitigation of random noise through averaging over multiple readouts and the extraction of information from the nonlinear dynamics induced by $\Omega_{i}S_{x,y}^{(i)}\sim\sum_{j\neq i}S_{z}^{(j)}S_{x,y}^{(i)}$.
.6 VI.    Weather forecasting and echo state networks
The weather dataset referenced in the main text is sourced from Kaggle dai (2019), a prominent online platform for data science competitions and resources. The dataset provides daily weather data for Delhi, India, covering 1,575 days from January 1, 2013, to April 24, 2017. It includes four features: temperature, humidity, wind speed, and air pressure. For simplicity of experiments, this study focuses on the first two features, temperature and humidity, whose values are presented in FIG. 8. Direct multi-step forecast is performed for the two-dimensional temporal sequence by training separate reservoir weights to model the mapping from $\{{\bf y}_{k}\}$ to $\{{\bf y}_{k+h}\}$ for different forecast horizons $h$, effectively enabling multitasking Taieb et al. (2012). The dataset is partitioned into 374 steps for washout, 600 steps for training, and the remaining steps for testing. For the testing phase, 600 steps are used for one-step-ahead prediction, as shown in Fig. 4(a) of the main text, while 556 steps are used for multi-step-ahead prediction with a maximum forecast horizon of $h_{\max}=45$.
Figure 8: Temperature and humidity trajectories from the daily weather dataset. The vertical dashed lines indicate the boundaries between the washout, training, and testing sets.
Figure 9: Multi-step-ahead forecasting performance results obtained using support vector regression with a radial basis function kernel as a nonlinear processing method on the reservoir readouts.
The baseline model employs Echo State Networks (ESNs), a widely used reservoir computing model based on recurrent neural networks Jaeger and Haas (2004). For an ESN with $M$ network nodes (denoted as ${\rm ESN}(M)$), the state update and output are described by
${\bf x}_{k}=\tanh\left({\bf W}_{\rm in}\begin{bmatrix}1\\ {\bf s}_{k}\end{bmatrix}+{\bf W}{\bf x}_{k-1}\right),~~y_{k}=\begin{bmatrix}1&{\bf x}^{\top}_{k}\end{bmatrix}{\bf w}_{\rm out},$ (19)
where ${\bf W}_{\rm in}\in\mathbb{R}^{M\times(1+d_{\rm in})}$, ${\bf W}\in\mathbb{R}^{M\times M}$ and ${\bf w}_{\rm out}$ are the input, recurrent, and output weights, respectively. Only the output weights ${\bf w}_{\rm out}$ are trained to minimize the prediction error, while ${\bf W}_{\rm in}$ and $\bf W$} are randomly initialized beforehand and remain fixed during training.
We implemented the ESNs using the Python package reservoirpy Trouvain et al. (2020), where ${\bf W}_{\rm in}$ is sampled from a Bernoulli distribution over $\{-1,1\}$ and $\bf W$} is sampled from a standard normal distribution. The connectivity $k$ and spectral radius $r$ of of the recurrent weight matrix $\bf W$}, defined as the fraction of nonzero entries and the largest absolute eigenvalue, respectively, are closely related to the computational power of ESNs. In this work, these hyperparameters were determined via grid search, yielding $(k,r)$ pairs of (0.025, 0.99), (0.01, 0.99), (0.0025, 0.99), and (0.0025, 0.99) for ESN(500), ESN(1000), ESN(5000), and ESN(10000), respectively. FIG. 11 presents the 45-day-ahead prediction results of different models.
We evaluated support vector regression (SVR) with a radial basis function (RBF) kernel as a nonlinear post–processor for the reservoir readouts. Our implementation employed the SVR module from the scikit-learn library Pedregosa et al. (2011), with the regularization parameter $C$ and the tolerance margin $\epsilon$ systematically tuned. Despite exploring a wide range of hyperparameter settings, SVR did not deliver significant performance improvements for ESNs. FIG. 9 reports the best results obtained.
.7 VII.    Effects of $T_{1}$ and $T_{2}$ relaxation on the fading memory in quantum reservoir computing
In this section, we examine the influence of the $T_{1}$ and $T_{2}$ relaxation on the fading memory of the quantum reservoir, as quantified by the short-term memory (STM) capacity, through numerical simulations. STM capacity quantifies the ability of the reservoir to reconstruct past inputs from its current state, which is a fundamental property for processing temporal information.
First, we reiterate the necessity of incorporating relaxation processes. As mentioned in the main text, an effective quantum reservoir should possess a high-dimensional state space, nonlinear input–output transformations, and fading memory. In unitary systems (commonly employed in quantum algorithm frameworks) the first two properties emerge naturally: the exponential scaling of the Hilbert space provides a high-dimensional representation, while nonlinearity arises from two mechanisms: (i) the input encoding (e.g., rotations yielding sine and cosine responses) Govia et al. (2022) and (ii) the multiplicative interaction between sequential inputs Fujii and Nakajima (2017); Pérez-Salinas et al. (2020). In contrast, fading memory requires that, after a sufficiently long input sequence, the reservoir state converges to a unique trajectory independent of its initial conditions. Due to the distance-preserving nature of unitary dynamics, nonunitary processes (such as $T_{1}$ and $T_{2}$ relaxation) are essential for gradually erasing old information.
We then assess how these relaxation processes affect the STM capacity. For a random input sequence $\{s_{k}\}$, the target output for a delay $t_{d}$ is defined as
$y_{k}=s_{k-t_{d}}.$ (20)
The STM capacity for delay $t_{d}$, denoted by $C(t_{d})$, is given by
$C(t_{d})=\frac{{\rm cov}^{2}({\bf y},{\bf\hat{y}})}{\sigma^{2}({\bf y})\sigma^{2}({\bf\hat{y}})}$, (21)
where $\bf y$} and $^{\bf\hat{y}}$ represent the target and predicted output sequences, respectively. The total STM capacity is then obtained by summing over all delays,
$C_{\rm STM}=\sum_{t_{d}=0}^{t_{d}^{\max}}C(t_{d}),$ (22)
where $t_{d}^{\max}$ is chosen to 100 in our simulations such that $C(t_{d})$ becomes negligible for larger delays.
Freq.(Hz) C H F $T_{1}$(s) $T_{2}$(ms)
C 79.8 2.9 203
H 160.6 201.0 2.8 236
F -194.3 48.0 157.2 3.1 200
Table 3: Parameters of the three-spin diethyl fluoromalonate sample. The left portion presents chemical shifts (diagonal elements) and $J$-couplings (off-diagonal elements), while the right lists the corresponding relaxation times.
In our numerical simulations, we employ a three‑spin system with parameters drawn from diethyl fluoromalonate, a widely used NMR molecular sample, as detailed in Table 3. The spin polarizations are set to $1,\,3.98,\,3.74]\times 10^{-4}$ for carbon, proton, and fluorine respectively. Considering that the input interval affects the memory capacity, evolution duration $\tau$ is varied from $10^{-4}~{\rm s}$ to $10^{1}~{\rm s}$ to explore its influence. To avoid the uncertainty caused by the selection of partial measurements, the readout is obtained by measuring $4^{3}-1=63$ Pauli operators to capture the full state information. These expectation values are then recalibrated by dividing by the polarization and further perturbed by adding Gaussian noise with $\sigma=10^{-5}$ to improve the robustness of linear regression. For the dataset, we generate a prolonged random sequence $\{s_{k}\}$ uniformly distributed in $[0,1]$. The training and test sets are taken from the tail of the sequence, with sizes of 3000 and 1000 points, respectively. A washout period of size $\max\left\{\left\lceil 10/\tau\right\rceil,t_{d}^{\max}\right\}$ is applied immediately prior to the training data to eliminate the effects of the initial state, lasting approximately 10 seconds.
Figure 10: (a) Total short-term memory (STM) capacity $C_{\rm STM}$ as a function of the input interval $\tau$, comparing four scenarios: amplitude damping only (AD, corresponding to $T_{1}$ relaxation), phase damping only (PD, corresponding to $T_{2}$ relaxation), combined amplitude and phase damping (AD & PD), and unitary dynamics (U). (b) and (c) depict the STM capacity $C(t_{d})$ versus delay $t_{d}$ for the combined amplitude and phase damping case and for amplitude damping only, respectively.
FIG. 10(a) illustrates the total STM capacity $C_{\rm STM}$ for various configurations of input intervals and relaxation mechanisms. As expected, $C_{\rm STM}$ vanishes under unitary evolution due to the absence of fading memory. In a unitary system, the reservoir state retains the full influence of the initial state and all prior inputs, preventing the selective extraction of relevant recent information. However, although the introduction of $T_{2}$ relaxation (phase damping) leads to some information erasure, it alone is insufficient to achieve a meaningful fading memory. Under unitary dynamics, the quantum system continuously interconverts population and coherence, while dephasing gradually eliminates coherence. This interplay ultimately depolarizes the reservoir, driving it toward the trivial maximally mixed state $\mathds{1}/2^{N}$.
Thus, not every dissipative process is suitable for quantum reservoir computing; the key requirement is a process that drives the system toward a nontrivial state—a role fulfilled by $T_{1}$ relaxation (generalized amplitude damping). Several studies have underscored the necessity of non-unital channels (i.e., channels that do not preserve the identity) in quantum reservoir computing Mart´ınez-Peña and Ortega (2023); Monzani et al. (2024). From an information-theoretic perspective, we conjecture that entropy-decreasing processes are essential for achieving effective fading memory. In amplitude damping channels, the system is driven toward an equilibrium state with lower entropy than the maximally mixed state. Moreover, previous schemes employing input qubit reset Fujii and Nakajima (2017) and probabilistic state reset Chen et al. (2020) incorporate entropy-reducing state preparation processes. In contrast, unitary operations preserve entropy, whereas dephasing increases it Nielsen and Chuang (2001).
The configuration incorporating $T_{1}$ relaxation exhibits a nonzero STM capacity, with $C_{\rm STM}$ initially increasing and then decreasing as the input interval varies. FIG. 10(b) and (c) show the corresponding behavior of $C(t_{d})$ as a function of delay. In the long-time limit ($\tau\gg T_{1}$), the reservoir relaxes to an equilibrium state independent of the inputs, causing $C_{\rm STM}$ to vanish. For relatively large $\tau$, relaxation erases most of the input history while only retains the most recent inputs, manifested as a narrow plateau ($C(t_{d})\approx 1$) followed by a rapid decline in $C(t_{d})$. As $\tau$ decreases, the reservoir incorporates information from more input steps; however, nonlinear mixing of these inputs can reduce the effectiveness of linear memory retrieval, resulting in an initial $C(0)<1$ and a gradual decay to zero at larger $t_{d}$. With further decrease in $\tau$, the overall values of $C(t_{d})$ diminish, and in the short-$\tau$ limit, the system behaves similarly to a unitary system, lacking effective fading memory. These results suggest the existence of an optimal input interval that balances the number of retained inputs with high memory fidelity at each delay.
The results incorporating both $T_{1}$ and $T_{2}$ relaxation (combined amplitude and phase damping) indicate that even in the presence of strong decoherence, the quantum reservoir remains capable of nontrivial information processing. In the case of only amplitude damping, the effective dephasing rate is reduced to $(2T_{1})^{-1}$. Consequently, the optimal total STM capacity is reached at larger $\tau$ compared to the combined $T_{1}$ and $T_{2}$ case, as $T_{1}$ is roughly one order of magnitude larger than $T_{2}$ (see TABLE 3). Although one might expect that omitting $T_{2}$ dephasing would better preserve information encoded in coherence, simulations show no significant improvement in STM capacity. This is likely because a slow relaxation ($\sim 10^{-1}~{\rm Hz}$) tends to take effect over a long evolution ($\sim 10^{0}~{\rm s}$), during which the high-frequency Hamiltonian ($\sim{10^{2}}~{\rm Hz}$) causes the inputs to be largely mixed in a nonlinear manner. In this context, the additional dephasing appears beneficial for managing memory. Overall, improved performance is anticipated by jointly tuning both the relaxation rate and the input interval to suit specific task requirements Sannia et al. (2024).
Figure 11: 45-day-ahead predictions of temperature and humidity. The blue and orange curves represent the target and predicted trajectories, respectively. From top to bottom, the models used are: ESN(500), ESN(1000), ESN(5000), ESN(10000), QRC with linear regression, and QRC with support vector regression using a nonlinear radial basis function kernel. The predictions of ESNs are averaged over 100 realizations, with shaded regions representing the standard deviation.